### v1

@@ Abstract
Robot navigation is increasingly crucial across applications like delivery services and warehouse management. The integration of Reinforcement Learning (RL) with classical planning has given rise to meta-planners that combine the adaptability of RL with the explainable decision-making of classical planners. However, the exploration capabilities of RL based meta-planners during training are often constrained by the capabilities of the underlying classical planners. This constraint can result in limited exploration, thereby leading to sampling skew issues. To address these issues, our paper introduces a novel framework, Digimon, which begins with behavior-guided diagnosis for exploration bottlenecks within the meta-planner and follows up with a mitigation strategy that conducts up-sampling from diagnosed bottleneck data. Our evaluation shows 13.5%+ improvement in navigation performance, greater robustness in out-of-distribution environments, and a 4$\times$ boost in training efficiency. Digimon is designed as a versatile, plug-and-play solution, allowing seamless integration into various RL-based meta-planners.
@@ I Introduction
Robot navigation is essential for applications like delivery services, hospital logistics, warehouse management, and library automation. Navigating cluttered environments remains a key challenge for achieving agile, efficient, and safe movement. Classical planners [abdulsaheb2022robot, asano1985visibility, Canny1985AVM, Sampling2014] address this by using specific cost functions and rules, which are difficult to generalize and often require expert manual tuning [Campbell2020, raheem2022robot, zheng2021rostuning, feng2024rocas]. In contrast, learning-based approaches [pfeiffer2017perception, gao2017intention, chiang2019learning, xiao2021toward, liu2021lifelong, xiao2021agile, xiao2020appld] offer adaptability but can suffer from instability and lack interpretability, leading to erratic behaviors that compromise safety and efficiency.
To overcome these limitations, recent studies introduced meta-planners [xu2021applr, xiao2022motion] that integrate RL with classical planning frameworks (see Fig. 1). RL-based meta-planners dynamically adjust the parameters of classical planners, combining RL’s adaptability with the explainable decision-making of classical methods for balanced robot navigation.
However, existing meta-planners suffer from sampling skew due to the integration of RL and classical planners. The exploration capabilities of RL are often limited by the classical planner’s inherent strategies. For instance, if the classical planner prefers conservative paths, the RL meta-planner may be unable to explore more aggressive or unconventional routes, resulting in sub-optimal performance. Previous approaches [swazinna2021overcoming, shen2023loose, mark2024offline, Strehl2005ATA, nikishin2022primacy], such as adding noise to RL actions, do not address these specific integration challenges. Therefore, there is a crucial need for specialized methods to mitigate the biases introduced by combining RL meta-planners with classical planning.
@@
(b) RL End-to-End
(c) Hybrid
Figure 1: Different Types of Planners. Note that the meta-planner is applied onto classical local planner $f$, not global trajectory planner.
In this paper, we propose Digimon, a diagnosis and mitigation framework addressing the limitations of RL meta-planners. Digimon starts with a behavior-guided diagnostic phase to identify exploration bottlenecks caused by classical planners. It then employs a novel up-sampling strategy to sample data from these bottleneck areas, providing RL meta-planner with a balanced replay buffer. This enhances RL exploration ability and mitigates sampling skew caused by classical planner’s preference. In summary, our contributions are:
  • We identify and analyze the exploration bottlenecks and sampling skews arising from integrating RL with classical planning in meta-planners.
  • We present Digimon, featuring a diagnostic phase to detect bottlenecks and a mitigation phase that uses data up-sampling to improve RL meta-planner training.
  • Digimon achieves 13.5%+ improvement in navigation performance, greater robustness in out-of-distribution environments, and a 4$\times$ increase in training efficiency. Additionally, Digimon is a versatile, plug-and-play solution, compatible with various RL-based meta-planners.
@@ II-A Motion Planning
Motion planning is used to determine a trajectory for a robot to move from a start position to a goal position while avoiding obstacles (as visualized in Fig. 2). It is worth noting that there are 2 major categories of planning algorithms: local trajectory planners (e.g. DWA [fox1997dwa], TEB [teb]) and global path planners (e.g., Dijkstras [dijkstra], A* [astar], RRT [lavalle1998rrt, lavalle2001rrt]).
@@
II-B Meta-Planner
Figure 2: Maps of various difficulty levels.
Meta-Planner operates on top of the original planner $f$, adjusting its configuration $\theta$. Fig. 1c shows that a hybrid planner with a meta-planner combines the explainability of classical planners with the adaptability of RL-based planners, enhancing performance. It has been applied across domains such as robot navigation [fridovich2018planning, xiao2022motion, siva2019robot, pokle2019deep], autonomous driving [jiang2022efficient], drone control [loquercio2022autotune], and robotic manipulation [Moll2021HyperPlanAF, jia2023dynamic].
There are two types of meta-planners: static and dynamic. Static meta-planners tune the configuration $\theta$ before deployment, while dynamic meta-planners adjust it in real-time.
@@ II-B Meta-Planner
Dynamic. Dynamic meta-planners adjust parameters on-the-fly for different scenarios [xiao2021learning, Binch2020, xiao2020appld, xu2021applr]. For example, [xiao2020appld] learns a library of parameter sets and switches based on context, while [xu2021applr] uses reinforcement learning to optimize parameter policies over decision sequences.
II-C RL as Meta-Planner
While existing works [gu2022constrainedRL, gehring2022rl4classical] combine RL with rule-based classical planning to enhance global path planning, our paper focuses on integrating RL with a local planner.
We formulate the navigation problem with a meta-planner as a Markov Decision Process (MDP) $\mathcal{M}$, i.e., a tuple $(\mathcal{S},\mathcal{A},\mathcal{T},\gamma,\mathcal{R})$. $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is the state transition function, $\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow R$ is the reward function, and $\gamma$ is the discount factor. At each time step t, the RL agent takes the current state $s_{t}\in\mathcal{S}$ as input and outputs an action $a_{t}\in\mathcal{A}$. Then the environment will transit to the next state $s_{t+1}\sim\mathcal{T}(\cdot|s_{t},a_{t})$ and the agent will receive a reward $r_{t}=\mathcal{R}(s_{t},a_{t})$. The overall objective of RL agent is to learn a policy $\pi:\mathcal{S}\rightarrow A$ that can be used to select actions that maximize the expected cumulative reward over time, i.e. $J=\mathbb{E}_{(s_{t},a_{t})\sim\pi}[\sum_{t=0}^{\infty}\gamma^{t}r_{t}].$
Within the MDP, we optimize a meta-planner policy that interacts with a composite environment consisting of the navigation world $\mathcal{W}$ (filled with obstacles) and a configurable motion planner $f$ parameterized by $\theta$ and receiving sensory inputs $o$. The state space, action space, reward function, and termination conditions are defined as follows:
State space. Following [xu2021applr], we assume a local goal $g$ is always reachable and acts as a waypoint on a global path. The direction to $g$, $\phi\in[-\pi,\pi]$, relative to the robot’s orientation, provides directional information. Thus, the state is $s_{t}=(o_{t},\phi_{t},\theta_{t-1})$, where $o_{t}$ are sensory inputs, $\phi_{t}$ the angle to $g$, and $\theta_{t-1}$ the previous planner configuration.
Action space. The agent’s action $a_{t}=\pi(s_{t})$ updates the planner’s parameters $\theta_{t}$. Executing $a_{t}$ applies $\theta_{t}$ to planner $f$, moving the robot and transitioning to the next state $s_{t+1}$.
Reward. The agent receives an immediate reward $r_{t}$ inversely related to its distance to $g$. An episode ends when a maximum time $T$ is reached or a termination condition is triggered, with the final reward penalizing timeouts and collisions.
The goal of the RL agent within the meta-planner is to learn a policy $\pi$ that can select a planner parameter $\theta_{t}$ for $f$ to achieve optimal navigation performance. The objective can be formulated as:
$\max_{\begin{subarray}{c}\pi\end{subarray}}J^{\pi}=\mathbb{E}_{s_{0},\theta_{t}\sim\pi(s_{t}),s_{t+1}\sim\mathcal{T}(s_{t},\theta_{t})}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}r_{t}\Big{]}.$ (1)
II-D Prioritized Experience Replay
Prioritized Experience Replay [schaul15prioritized] improves RL by assigning higher probabilities to more informative experiences in the replay buffer, enhancing learning efficiency. It selects samples that significantly contribute to the agent’s progress. In contrast, our approach focuses on improving data collection quality by reducing sampling bias and promoting better exploration, resulting in a more informative replay buffer. Together, these methods complement each other to advance RL efficiency.
II-E Roadmap Analysis
@@
III Sampling Skew of Existing Methods
APPLR [xu2021applr], the state-of-the-art dynamic meta-planner, exhibits a sampling skew that can lead to inefficiencies and limit performance improvements. In this section, we discuss the sampling skew, key insights, and related technical challenges.
@@ III Sampling Skew of Existing Methods
Figure 3: Sampling skew in existing RL-based Meta-Planner. For convenient visualization, we show 2D projections of the difficult map from Fig. 2 and omit the robot body. The start position is on the left of the map and the goal is on the right. Red cross markers are the robot trajectories at each timestamp. On top and right of the maps are the frequency distribution of robot positions $x$ and $y$.
@@
Meta-Planner Training. Fig. 3 shows 2D projections of the difficult map from Fig. 2, where the robot navigates obstacles from the start (left) to the goal (right) without collisions and within a time limit. The meta-planner adjusts planner’s configurations, i.e. actions $a_{t}=\pi(s_{t})$, at a fixed frequency (e.g., 1 Hz) in response to environmental changes, and we record the robot’s positions for reward calculation and policy training.
@@
Fig. 3(a) shows the early-stage challenges with a randomly initialized policy, where the robot often gets stuck (Area A). Over time (Fig. 3(b)), the policy improves, avoiding early obstacles but still struggling with tight spaces. As training progresses (Fig. 3(c)), the robot reduces getting stuck but takes inefficient detours (Area C), leading to timeout. Further training (Fig. 3(d)) improves navigation, but the robot still faces difficulties passing through narrow gates (Area D).
Sampling Skew Phenomenon. From training iterations in Fig. 3, sampling skew can be summarized by 3 observations:
@@
Observation I: Challenging scenarios require more sampling. In difficult situations like dead corners (Fig. 3(b)), detours (Fig. 3(c)), and narrow gates (Figures 3(a), 3(d)), the classical planner’s limitations and insufficient sampling of the meta-planner often cause delays, timeouts, or collisions. These critical hard-to-pass areas require significantly more sampling to improve navigation efficiency and safety.
@@
Observation II: Challenging scenarios have impact on subsequent scenarios. Hard-to-pass areas make the following scenarios (i.e., hard-to-reach) more difficult to access. This is reflected in the robot’s position distribution, showing higher frequency near hard-to-pass areas (e.g., Area A in Fig. 3(a) and Area B in Fig. 3(b)) and lower frequency in subsequent regions.
@@
Observation III: Challenging scenarios tend to be under-fitted. Certain challenging scenarios remain under-fitted, risking mission success due to two key factors. For instance, Area D initially remains hard-to-reach (Fig. 3(a)-3(c)) and is not effectively trained. As training progresses (Fig. 3(d)), Area D becomes hard-to-pass, as the robot reaches it smoothly but takes around 20 steps (about 20 seconds) per trajectory. This lengthy process limits efficient training within a fixed budget, making it difficult to fully train Area D.
Challenges. A simple solution to mitigate sampling skew is to focus on training the meta-planner more in under-fitted areas, but there is no automatic way to identify these regions. Analyzing static maps alone is insufficient (e.g., detecting narrow parts or sharp turns), as this only considers static information and ignores how the robot dynamically executes and interacts with its environment.
Our key insight is that hard-to-reach areas stem from preceding hard-to-pass areas. If there are no hard-to-pass areas, the robot can easily and safely reach its goal. Therefore, our robot behavior-guided method focuses on: (1) increasing sampling in hard-to-pass areas based on Observations I and II, and (2) reducing the time needed to initiate such sampling, aligning with Observation III. In Section IV, we describe how we identify these areas to address issues in meta-planner.
IV Design of Digimon
We introduce Digimon, consisting of two components: diagnosis (Section IV-B) and mitigation (Section IV-C) of sampling skew phenomenon, as discussed in Section III.
@@ IV-A Overview
Alg. 1 illustrates the integration of Digimon within the meta-planner policy training pipeline. Components enhanced by Digimon are highlighted in blue, indicating the modifications to the standard meta-planner policy. The remaining parts abstract the conventional components of a reinforcement learning training pipeline.
@@
Algorithm 1 Digimon for Enhancing RL-based Meta-Planner Training
1:Input: Initial meta-planner policy network $\pi$, high-resistance area $\mathcal{H}$, environment $\mathcal{E}=f\times\mathcal{W}$ for $\pi$ (where $f$ is classical planner and $\mathcal{W}$ is physical map).
2:Constants: Robot’s initial position ${\tt INIT\_POSE}$, threshold $\lambda$ for sampling initial pose from $\mathcal{H}$, policy training iteration $N$, trajectory collection number $K$, policy update number $L$.
3:Output: The converged meta-planner policy network $\pi^{*}$.
@@
10:        if ${\tt getRandom}($0,1)))) < $\lambda$ and $\mathcal{H}\neq\emptyset$ then
@@
13:        for $t=0,...,T$ do
14:           Run $\pi$ in $\mathcal{E}$, take action $a_{t}=\pi(s_{t})$.
15:           Obtain reward $r_{t}$, state $s_{t+1}$ and pose $p_{t+1}$.
@@
17:           if $\pi$ gets ${\tt done}$ signal from $\mathcal{E}$ then,
@@
22:        $\mathcal{B}\leftarrow{\tt getMiniBatch}(\mathcal{D})$triangleright Random mini-batch
23:        Update policy $\pi$ using $\mathcal{B}$. $\triangleright$ Algorithm-specific
24:Return the final meta-planner policy $\pi^{*}$.
IV-B Sampling Skew Diagnosis

### v3

@@ Abstract
Robot navigation systems are critical for various real-world applications such as delivery services, hospital logistics, and warehouse management. Although classical navigation methods provide interpretability, they rely heavily on expert manual tuning, limiting their adaptability. Conversely, purely learning-based methods offer adaptability but often lead to instability and erratic robot behaviors. Recently introduced parameter tuners aim to balance these approaches by integrating data-driven adaptability into classical navigation frameworks. However, the parameter tuning process currently suffers from training inefficiencies and redundant sampling, with critical regions in environment often underrepresented in training data.
In this paper, we propose EffiTune, a novel framework designed to diagnose and mitigate training inefficiency for parameter tuners in robot navigation systems. EffiTune first performs robot-behavior-guided diagnostics to pinpoint critical bottlenecks and underrepresented regions. It then employs a targeted up-sampling strategy to enrich the training dataset with critical samples, significantly reducing redundancy and enhancing training efficiency. Our comprehensive evaluation demonstrates that EffiTune achieves more than a 13.5% improvement in navigation performance, enhanced robustness in out-of-distribution scenarios, and a 4× improvement in training efficiency within the same computational budget.
@@ I Introduction
Robot navigation is essential for applications like delivery services, hospital logistics, warehouse management, and library automation. Navigating cluttered environments remains a key challenge for achieving agile, efficient, and safe movement. Classical navigation systems [abdulsaheb2022robot, asano1985visibility, Canny1985AVM, Sampling2014] address this by using specific cost functions and rules, which are difficult to generalize and often require expert manual tuning [Campbell2020, raheem2022robot, zheng2021rostuning, feng2024rocas, kate2025roscallbax]. In contrast, learning-based approaches [pfeiffer2017perception, gao2017intention, chiang2019learning, xiao2021toward, liu2021lifelong, xiao2021agile, xiao2020appld] offer adaptability but can suffer from instability and lack interpretability [feng2023decree, chen2024temporalrl, cheng2024fusion, cheng2024badpart], leading to erratic behaviors that compromise safety and efficiency.
To overcome these limitations, recent studies introduced parameter tuner [xu2021applr, xiao2022motion, jiang2022efficient, jiang2023learning] that integrate data-driven tuner with classical planning frameworks (see Fig. 1). These data-driven parameter tuners adjust the parameters (across perception, planning and control components) of classical navigation systems in different scenarios, combining adaptability of data-driven methods with the explainable decision-making of classical methods for balanced robot navigation.
However, existing parameter tuners encounter significant challenges. The interaction between classical navigation systems and the environment is computationally expensive, often leading to sampling inefficiencies. Furthermore, the training process of these tuners tends to include redundant sampling. In addition, our analysis reveals that only a few key regions in the environment are critical for achieving optimal performance. Unfortunately, these pivotal areas are underrepresented in the training data due to both the high cost of environmental interaction and the prevalence of redundant sampling. To address these challenges, our approach emphasizes the most informative data during training, significantly enhancing the effectiveness and efficiency of the parameter tuner. Unlike static roadmap analysis, which offers a fixed representation, our method captures the robot’s dynamic behavior, providing a more accurate and informative analysis.
In this paper, we propose EffiTune, a diagnosis and mitigation framework that addresses inefficiency in parameter tuner training. EffiTune begins with a robot-behavior-guided diagnostic phase that identifies critical bottlenecks. It then employs a novel up-sampling strategy to selectively enrich the training data with samples from these key regions, providing the parameter tuner with a more balanced and informative dataset while reducing redundancy. This targeted approach accelerates the training process of tuners, enhances adaptability, and ultimately improves overall navigation performance.
@@
(b) Hybrid
Figure 1: Different Types of Robot Navigation Systems. Parameter tuners are typically data-driven and requires training.
In summary, our contributions are:
  • We identify training inefficiency issue for parameter tuner training in robot navigation.
  • We introduce EffiTune, a framework that features a diagnostic phase to detect these critical regions and a mitigation phase that uses targeted data up-sampling to accelerate parameter tuner training.
  • EffiTune achieves a 13.5%+ improvement in navigation performance, greater robustness in out-of-distribution environments, and a 4× increase in training efficiency within the same timing budget.
@@ II-A Motion Planning
Motion planning is used to determine a trajectory for a robot to move from a start position to a goal position while avoiding obstacles (as visualized in Fig. 2). There are 2 major categories of planning algorithms: local trajectory planners[fox1997dwa, teb] and global path planners [dijkstra, astar, lavalle1998rrt, lavalle2001rrt].
@@
(a) 0 $\sim$ 2.5 s/m
(b) 2.5 $\sim$ 4.0 s/m
(c) 4.0 $\sim$ 8.0 s/m
Figure 2: Maps of various difficulty levels. We use the difficulty metric of “normalized traversal time” (traversal time averaged over the ten trials and normalized by path length) from [perille2020benchmarking] to partition map datasets to 3 levels.
II-B Parameter Tuner
Parameter Tuner $\tau$ operates on top of the original planner $f$, taking environmental states $s$ as input and adjusting navigation configuration $\theta$. Fig. 1c shows that a hybrid planner with a parameter tuner combines the explainability of classical planners with the adaptability of data-driven (e.g., end-to-end) planners, enhancing performance. It has been applied across domains such as robot navigation [fridovich2018planning, xiao2022motion, siva2019robot, pokle2019deep], self-driving [jiang2022efficient], autonomous system [feng2025intentest, zheng2025llm], drone control [loquercio2022autotune], and robotic manipulation [Moll2021HyperPlanAF, jia2023dynamic].
There are two types of parameter tuners: static and dynamic. Static parameter tuners tune the configuration $\theta$ before deployment, while dynamic parameter tuners adjust it in real-time.
@@ II-B Parameter Tuner
Dynamic. Dynamic parameter tuners adjust parameters on-the-fly for different scenarios [xiao2021learning, Binch2020, xiao2020appld, xu2021applr]. For example, [xiao2020appld] learns a library of parameter sets and switches based on context, while [xu2021applr] uses deep neural network to optimize parameters under various situations.
II-C Roadmap Analysis
@@
III Training Inefficiency of Existing Methods
APPLR [xu2021applr], the state-of-the-art dynamic parameter tuner, exhibits a training inefficiency that can lead to limited performance improvements. In this section, we discuss the training inefficiency, key insights, and related technical challenges.
@@ III Training Inefficiency of Existing Methods
Figure 3: Training inefficiency in existing parameter tuner. For convenient visualization, we show 2D projections of the difficult map from Fig. 2 and omit the robot body. The yellow arrow denotes the start position and the yellow circle denotes the goal. Red cross markers are the robot trajectories at each timestamp (each subfigure contains 3 trajectories to show the distribution). On top and right of the maps are the frequency distribution of robot positions $x$ and $y$.
@@
Parameter Tuner Training. Fig. 3 shows 2D projections of the difficult map from Fig. 2, where the robot navigates obstacles from the start (left) to the goal (right) without collisions and within a time limit. The parameter tuner adjusts planner’s configurations, i.e. adjustment $a_{t}=\tau(s_{t})$, at a fixed frequency (e.g., 1 Hz) in response to environmental changes, and we record the robot’s positions for reward calculation and tuner training.
@@
Fig. 3(a) shows the early-stage challenges with a randomly initialized tuner, where the robot often gets stuck (Area A). Over time (Fig. 3(b)), the tuner improves, avoiding early obstacles but still struggling with tight spaces. As training progresses (Fig. 3(c)), the robot reduces getting stuck but takes inefficient detours (Area C), leading to timeout. Further training (Fig. 3(d)) improves navigation, but the robot still faces difficulties passing through narrow gates (Area D).
Training Inefficiency. From training iterations in Fig. 3, training inefficiency can be summarized by 3 observations:
@@
Observation I: Challenging scenarios require more sampling. In situations like dead corners (Fig. 3(b)), detours (Fig. 3(c)), and narrow gates (Figures 3(a), 3(d)), the classical planner’s limitations and insufficient sampling of the parameter tuner often cause delays, timeouts, or collisions. These critical hard-to-pass areas require significantly more sampling to improve navigation efficiency and safety.
@@
Observation II: Challenging scenarios prevent training on subsequent scenarios. Hard-to-pass areas make the following scenarios (i.e., hard-to-reach) more difficult to access. This is reflected in the robot’s position distribution, showing higher frequency near hard-to-pass areas (e.g., Area A in Fig. 3(a) and Area B in Fig. 3(b)) and lower frequency in subsequent regions.
@@
Observation III: It is hard and expensive to train on challenging scenarios. Certain challenging scenarios remain under-fitted, risking mission success due to two key factors. For instance, Area D initially remains hard-to-reach (Fig. 3(a)-3(c)) and is not accessible to the robot. As training progresses (Fig. 3(d)), Area D becomes accessible, as the robot can reach it smoothly but takes around 20 steps (about 20 seconds) per trajectory to arrive in Area D. This lengthy process is unnecessary but expensive and makes the tuner difficult to sufficiently train on Area D, leading to inefficient training within the given time budget.
Our Insight. Observations I and II highlight the importance of accurately identifying challenging regions and adequately training within them. However, Observation III reveals that extensive training in these areas is computationally costly, necessitating efficiency improvements.
To address these issues, we propose a two-stage approach: (1) leveraging a robot-behavior-guided method to pinpoint critical and challenging regions effectively, and (2) employing targeted up-sampling strategies to reduce the overhead associated with sampling initiation. In Section IV, we detail our methodology for identifying these critical regions, directly addressing the inefficiency issue in parameter tuner training.
IV Design of EffiTune
We introduce EffiTune (shown in Fig. 4), with two stages: training inefficiency diagnosis (Section IV-B) and mitigation (Section IV-C) for parameter tuners.
@@ IV-A Overview
Alg. 1 illustrates the integration of EffiTune within the parameter tuner training pipeline. Components enhanced by EffiTune are highlighted in blue, indicating the modifications to the standard parameter tuner. The remaining parts abstract the conventional components of a deep learning training pipeline.
@@
Figure 4: Overview of EffiTune Framework
Algorithm 1 EffiTune for Accelerating Parameter Tuner Training
1:Input: Initial parameter tuner $\tau$, high-resistance area $\mathcal{H}$, environment $\mathcal{E}=f\times\mathcal{W}$ for $\tau$ (where $f$ is classical planner and $\mathcal{W}$ is physical map).
2:Constants: Robot’s initial position $\tt INIT\_POSE$}, threshold $\lambda$ for sampling initial pose from $\mathcal{H}$, tuner training iteration NN, trajectory collection number $K$, tuner update number $L$.
3:Output: The converged parameter tuner $\tau^{*}$.
@@
10:   if ${\tt getRandom}($0,1)) < $\lambda$ and $\mathcal{H}\neq\emptyset$ then
@@
13:   for t=0,,Tt=0,...,T do
14:     Run $\tau$ in $\mathcal{E}$, take tuning adjustment $a_{t}=\tau(s_{t})$.
15:     Obtain feedback $r_{t}$, state $s_{t+1}$ and pose $p_{t+1}$.
@@
17:     if $\tau$ gets $\tt done$} signal from $\mathcal{E}$ then,
@@
22:   $\mathcal{B}\leftarrow{\tt getMiniBatch}(\mathcal{D})$ $\triangleright$ Random mini-batch
23:   Update tuner $\tau$ using $\mathcal{B}$. $\triangleright$ Algorithm-specific
24:Return the final parameter tuner $\tau^{*}$.
IV-B Training Inefficiency Diagnosis